Use Chebyshev’s theorem to find what percent of the values will fall between 175 and 241 for a data set with a mean of 208 and standard deviation of 11.
Use the Empirical Rule to find what two values 99.7% of the data will fall between for a data set with a mean of 245 and standard deviation of 16.


Answer by oscargut(2103)

Using Chebyshev’s theorem
P(175= 1-1/9 = 8/9
Answer:88.9%
Using empirical rule: values are: 208-3(11)= 175 and 208+3(11)=241
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com